[Brahms-l] Shift report 20040117 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Sat Jan 17 2004 - 15:59:19 EST
------------------------------------------------------------
Date: Sat Jan 17 15:59:18 2004, Shift: 08:00-16:00
Supervisor: rw, with: 
Log book#15 pages 35 - 35

Shift Summary:
------------------------------------------------------------
1. Beam from the previous shift is ended at 11:10am.

2. run No      trg3      trg6
   9725         ~10k       ~30k     (lost event numbers by mistake)
   9726    TPC Pedestal run
   9727    Non-TPC Pedestal run
   9732         20k        60k      End of both FS, MRS magnet setings

3. Some of DCs had high current readings, but after On/OFF operation, all
   becomes normal except T3A21(No HV),T5A2(20.6/43.4/1.7/0.7)
   Some of C1/C4 HVs have orange ramps.
   (~25V off from the demanded voltages) 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jan 17 15:59:39 2004

This archive was generated by hypermail 2.1.8 : Sat Jan 17 2004 - 15:59:58 EST